# Suppressing unknown disturbances to dynamical systems using machine learning

### v4

@@ Abstract
Identifying and suppressing unknown disturbances to dynamical systems is a problem with applications in many different fields. Here we present a model-free method to identify and suppress an unknown disturbance to an unknown system based only on previous observations of the system under the influence of a known forcing function. We find that, under very mild restrictions on the training function, our method is able to robustly identify and suppress a large class of unknown disturbances. We illustrate our scheme with the identification of both deterministic and stochastic unknown disturbances to an analog electric chaotic circuit and with a numerical example where a chaotic disturbance to the Lorenz system is identified and suppressed.
@@ Introduction
Identifying and suppressing an unknown disturbance to a dynamical system is a problem with many existing and potential applications in engineering nudell2013graph ; upadhyaya2015power ; mathew2016pmu ; chen2015disturbance ; ferreira2016survey ; lee2018data ; wang2019detection ; delabays2021locating ; delabays2022locating , ecology meurant2012ecology ; battisti2016introduction , fluid mechanics bewley1998optimal ; bewley2000general ; juillet2014experimental , and climate change verbesselt2010phenological . Traditional control theory disturbance identification and suppression techniques usually assume either an existing model for the dynamical system, linearity, or that the disturbance can be observed (for reviews of existing methods see, for example, Refs. canfield2003active ; chen2015disturbance ). In this Article we present a method for real-time disturbance identification and suppression that relies solely on observations of the dynamical system when forced with a known training forcing function. Our method is based on the application of machine-learning techniques to dynamical systems. Such techniques have found many applications, including the forecast of chaotic spatiotemporal pathak2018model and networked srinivasan2022parallel dynamics, estimation of dynamical invariants from data pathak2017using , control of chaos canaday2021model , network structure inference banerjee2019using , and prediction of extreme events pyragas2020using and crises in non-stationary dynamical systems kong2023reservoir ; patel2023using . For a review of other applications and techniques, see Refs. tanaka2019recent ; nakajima2021reservoir . In most of these previous works, a machine learning framework is trained to replicate the nonlinear dynamics of the system based on a sufficiently long time series of the dynamics.
@@ Results
We begin by outlining our method for identifying the unknown disturbance ${\bf g}(t)$. We will illustrate our technique using reservoir computing, a type of machine learning framework particularly suited for dynamical systems problems nakajima2021reservoir . In our implementation, we assume that we run the system in Eq. (2) during the “training” interval $[-T,0]$, and collect a time-series of the observed state vector $\{\hat{\bf x}(-T),\hat{\bf x}(-T+\Delta t),\dots,\hat{\bf x}(0)\}$. These variables are fed to the reservoir, a high-dimensional dynamical system with internal variables ${\bf r}\in\mathbb{R}^{M}$, where $M$ is the size of the reservoir. Here, following pathak2017using ; pathak2018model , we implement the reservoir as
${\bf r}(t+\Delta t)=\tanh[A{\bf r}(t)+W_{\text{in}}\hat{\bf x}(t)+1],$ (3)
where the $M\times M$ matrix $A$ is a sparse matrix representing the internal structure of the reservoir network and the $M\times N$ matrix $W_{\text{in}}$ is a fixed input matrix. The reservoir output $\bf u$ is constructed from the internal states as $\bf u=W_{\text{out}}{\bf r}$, where the $N\times M$ output matrix $W_{\text{out}}$ is chosen so that $\bf u$ approximates as best as possible the known training forcing function ${\bf f}(t)$. The optimization can be done by minimizing the cost function
@@
${\bf r}(t+\Delta t)=\tanh[A{\bf r}(t)+W_{\text{in}}{\bf x}(t)+1].$ (5)
@@
$\frac{dx}{dt}$ $=\sigma(y-x)+g_{x}(t),$ (6)
$\frac{dy}{dt}$ $=x(\rho-z)-y+g_{y}(t),$ (7)
$\frac{dz}{dt}$ $=xy-\beta z+g_{z}(t),$ (8)
@@ Methods
The values used for the resistors are chosen to produce the appropriate coefficients in the Lorenz equations of $\sigma=10$, $\beta=8/3$, and $\rho=28$. The integrating capacitors of $47$ nF were chosen to provide oscillations on the order of 30 Hz. Resistors had a component tolerance of 1%, while the capacitors have a 5% tolerance. The multiplication was done with an AD633 analog multiplier, while the integrating circuit was based on an TL082 operational amplifier. The output of this circuit is shown in Fig. 5(b). The characteristic butterfly shape is readily apparent, with each output swinging around 4 volts peak to peak (Vpp). This analog circuit represents the “undisturbed system” described in the main text. While it is constructed to obey approximately the (scaled) Lorenz equations, the component tolerances make it, for practical purposes, an unknown system from which we can measure the state variables $x$, $y$, and $z$. Electronic noise adds an additional complication not present in our numerical simulations.

### v5

@@ Abstract
Identifying and suppressing unknown disturbances to dynamical systems is a problem with applications in many different fields. Here we present a model-free method to identify and suppress an unknown disturbance to an unknown system based only on previous observations of the system under the influence of a known forcing function. We find that, under very mild restrictions on the training function, our method is able to robustly identify and suppress a large class of unknown disturbances. We illustrate our scheme with the identification of both deterministic and stochastic unknown disturbances to an analog electric chaotic circuit and with numerical examples where a chaotic disturbance to various chaotic dynamical systems is identified and suppressed.
@@ Introduction
Identifying and suppressing an unknown disturbance to a dynamical system is a problem with many existing and potential applications in engineering nudell2013graph ; upadhyaya2015power ; mathew2016pmu ; chen2015disturbance ; ferreira2016survey ; lee2018data ; wang2019detection ; delabays2021locating ; delabays2022locating , ecology meurant2012ecology ; battisti2016introduction , fluid mechanics bewley1998optimal ; bewley2000general ; juillet2014experimental , and climate change verbesselt2010phenological . Traditional control theory disturbance identification and suppression techniques usually assume either an existing model for the dynamical system, linearity, or that the disturbance can be observed (for reviews of existing methods see, for example, Refs. canfield2003active ; chen2015disturbance ). In this Article we present a method for real-time disturbance identification and suppression that relies solely on observations of the dynamical system when forced with a known training forcing function. Our method is based on the application of machine-learning techniques to dynamical systems. Such techniques have found many applications, including the forecast of chaotic spatiotemporal pathak2018model and networked srinivasan2022parallel dynamics, estimation of dynamical invariants from data pathak2017using , control of chaos canaday2021model , network structure inference banerjee2019using , and prediction of extreme events pyragas2020using and crises in non-stationary dynamical systems kong2023reservoir ; patel2023using . For a review of other applications and techniques, see Refs. tanaka2019recent ; nakajima2021reservoir ; carrol2019Chaos . In most of these previous works, a machine learning framework is trained to replicate the nonlinear dynamics of the system based on a sufficiently long time series of the dynamics.
@@ Results
We begin by outlining our method for identifying the unknown disturbance ${\bf g}(t)$. We will illustrate our technique using reservoir computing, a type of machine learning framework particularly suited for dynamical systems problems nakajima2021reservoir . In our implementation, we assume that we run the system in Eq. (2) during the “training” interval $[-T,0]$, and collect a time-series of the observed state vector $\{\hat{\bf x}(-T),\hat{\bf x}(-T+\Delta t),\dots,\hat{\bf x}(0)\}$. These variables are fed to the reservoir, a high-dimensional dynamical system with internal variables ${\bf r}\in\mathbb{R}^{M}$, where $M$ is the size of the reservoir. Here, following pathak2017using ; pathak2018model , we implement the reservoir as the map
${\bf r}(t+\Delta t)=\tanh[A{\bf r}(t)+W_{\text{in}}\hat{\bf x}(t)+\beta],$ (3)
where the $M\times M$ matrix $A$ is a sparse matrix representing the internal structure of the reservoir network and the $M\times N$ matrix $W_{\text{in}}$ is a fixed input matrix. Here we choose the bias parameter $\beta=1$, which for our purposes nearly optimizes results (see Methods). The reservoir output $\bf u$ is constructed from the internal states as $\bf u=W_{\text{out}}{\bf r}$, where the $N\times M$ output matrix $W_{\text{out}}$ is chosen so that $\bf u$ approximates as best as possible the known training forcing function ${\bf f}(t)$. The optimization can be done by minimizing the cost function
@@
${\bf r}(t+\Delta t)=\tanh[A{\bf r}(t)+W_{\text{in}}{\bf x}(t)+1],$ (5)
@@
$\frac{dx_{L}}{dt}$ $=\sigma(y_{L}-x_{L})+g_{x}(t),$ (6)
$\frac{dy_{L}}{dt}$ $=x_{L}(\rho-z_{L})-y+g_{y}(t),$ (7)
$\frac{dz_{L}}{dt}$ $=x_{L}y_{L}-\beta z_{L}+g_{z}(t),$ (8)
@@
Figure 3: Identifying unknown disturbances: Rössler and Lorenz 96 systems. (Top) Unknown and reconstructed disturbance functions $[g_{x}(t),g_{y}(t)]$ (black curve) and $[u_{x}(t),u_{y}(t)]$ (red curve) along with the training forcing functions $[f_{x}(t),f_{y}(t)]$ (blue curves). For each case the reservoir was trained with $[f_{x}(t),f_{y}(t)]=[\cos(0.05t),\sin(0.05t)]$. (Bottom) Time series for the unknown (solid black) and recovered (dashed red) disturbance functions in the second component, $g_{y}(t)$ and $u_{y}(t)$ (or $g_{x_{2}}(t)$ and $u_{x_{2}}(t)$).
Next we present some additional results demonstrate generalizability of our mechanism. First, we consider an inverted version of our first example, namely we consider system dynamics defined by the Rössler system [i.e., Eqs. (9)–(11)] that are disturbed by time series arising from the Lorenz system [i.e., Eqs. (6)–(8)]. Specifically, we set $[g_{x}(t),g_{y}(t)]^{T}=[x_{L}(t)/20,y_{L}(t)/20]^{T}$ and use sinusoidal forcing, as before, for training, namely $[f_{x}(t),f_{y}(t)]^{T}=[\cos(t/20),\sin(t/20)]^{T}$. Results for this inverted example are plotted in Figs. 3(a) and (c), and show good agreement between the unknown and recovered disturbances. Second, to illustrate the efficacy of the methodology in a higher-dimensional system, we consider the Lorenz 96 model Lorenz96 whose variables $x_{i}$ for $i=1,\dots,N$ evolve according to
$\frac{dx_{i}}{dt}$ $=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F.$ (14)
Here we choose the dimension $N=8$ and set $F=8$ to realize high-dimensional chaos. We train the system with known forcing in the first two variables as in the prior example, $[f_{x_{1}}(t),f_{x_{2}}(t)]^{T}=[\cos(t/20),\sin(t/20)]^{T}$, then use an unknown disturbance of $[g_{x_{1}}(t),g_{x_{2}}(t)]^{T}=[\cos(t/2),\sin(11t/20)]^{T}$. Due to the high dimensionality we use a resevoir of twice the size as in other examples, namely, $M=2000$. Results for this example are plotted in Figs. 3(b) and (d), and show good agreement between the unknown and recovered disturbances.
@@
Figure 6: Suppressing unknown disturbances: Quasi-periodic and stochastic disturbances. For quasi-periodic (top) and stochastic (bottom) disturbances, the effectiveness of delayed control. In (a) and (e) the disturbances applied to the Lorenz system. For each case, (b)–(d) and (f)–(h) the disturbed (black) and undisturbed (red) attractors when using control gains $\alpha=0$, $10$, and $100$.
Figure 7: Suppressing unknown disturbances: Rössler system. (a) The disturbances applied to the Rössler system. (b)–(d) the disturbed (black) and undisturbed (red) attractors when using control gains $\alpha=0$, $10$, and $100$.
Figure 8: Suppressing unknown disturbances: Lorenz 96 system. (a) The disturbances applied to the Lorenz 96 system. (b)–(d) the disturbed (black) and undisturbed (red) attractors when using control gains $\alpha=0$, $4$, and $20$.
@@
We also present some additional results that demonstrate the generalizability of the suppression mechanism in the face of different types of disturbances. In particular, while we keep the undisturbed system defined by the Lorenz system, we consider first quasi-periodic disturbances that are composed from mismatched sinusoids, namely $[g_{x},g_{y},g_{z}]^{T}=[200\cos(2t/5)\sin(2\pi t/5),50\cos(t/2)+50\sin(\pi t),0]^{T}$, and, second, stochastic disturbances $[g_{x},g_{y},g_{z}]^{T}=[x_{S}(t),y_{S}(t),0]^{T}$ as defined in Eqs. (12) and (13) but with a stronger stochasitc component, specifically $D=75$. We plot the results from these two cases under delayed control in Fig. 6 on the top and bottom, respectively. First, in panels (a) and (e) we plot the time series of the disturbances applied to the Lorenz system, depicting the disturbances to the $x$ and $y$ components in solid blue and dashed red, respectively. Next, in panels (b) and (f) we plot the disturbed attractor without control (i.e., using $\alpha=0$) in black and also plotting the undisturbed attractor in red for comparison. Then in panels (c) and (g) we plot the disturbed and undisturbed attractors for $\alpha=10$, then in panels (d) and (h) for $\alpha=100$. As we increase the control gain $\alpha$ we see the disturbed attractor begins to resemble more so the undisturbed attractor.
Lastly we consider control of disturbances to other systems, specifically the Rössler system and the Lorenz 96 system with $N=8$ and $F=8$. In both cases we train the systems with sinusoidal forcing, $[f_{x}(t),f_{y}(t)]^{T}=[\cos(0.05t),\sin(0.05t)]^{T}$. Next, we disturb the $x$ and $y$ components of the Rössler system with the $x$ and $y$ components of the Lorenz system, $[g_{x},g_{y}]^{T}=[2x_{L(t)/5,2y_{L}(t)/5}]^{T}$ and we disturb the first two components of the Lorenz 96 system with sinusoids with offset frequencies, $[g_{x_{1}}(t),g_{x_{2}}(t)]^{T}=[50\cos(t/2),50\sin(11t/20)]^{T}$. In Figs. 7 and 8 we plot the results for the Rössler system and the Lorenz 96 system, respectively, plotting in panel (a) the disturbances applied to each, then in panels (b)–(d) the disturbed (black) and undisturbed (red) attractors as the control gain is increased: for the Rössler system we use $\alpha=0$, $10$, and $100$ and for the Lorenz 96 system we use $\alpha=0$, $4$, and $20$. (Note that for the Lorenz 96 system we were able to suppress disturbances quite well with even smaller control gains, thus the smaller values of $\alpha$ used.) In both cases we see that as the control gain is increased the disturbed dynamics get closer to the undisturbed dynamics.
@@ Methods
Choice of the Bias Parameter, $\beta$
Figure 9: Effect of the bias parameter. As a function of the bias parameter $\beta$, the combined mean squared error (MSE) between the unknown disturbances, $g_{x}$ and $g_{y}$, and those recovered by the reservoir, $u_{x}$ and $u_{y}$. Unknown disturbances were presented at different magnitudes, scaled by $\mu$ as $[g_{x},g_{y},g_{z}]^{T}=[\mu x_{R},\mu y_{R},0]^{T}$.
[…]
@@
The values used for the resistors are chosen to produce the appropriate coefficients in the Lorenz equations of $\sigma=10$, $\beta=8/3$, and $\rho=28$. The integrating capacitors of $47$ nF were chosen to provide oscillations on the order of 30 Hz. Resistors had a component tolerance of 1%, while the capacitors have a 5% tolerance. The multiplication was done with an AD633 analog multiplier, which has an error of $2\%$ of full scale, while the integrating circuit was based on an TL082 operational amplifier. The output of this circuit is shown in Fig. 10(b). The characteristic butterfly shape is readily apparent, with each output swinging around 4 volts peak to peak (Vpp). This analog circuit represents the “undisturbed system” described in the main text. While it is constructed to obey approximately the (scaled) Lorenz equations, the component tolerances make it, for practical purposes, an unknown system from which we can measure the state variables $x$, $y$, and $z$. Electronic noise and uncertainty from the analog multipliers adds an additional complication not present in our numerical simulations.
@@
Data Availability
The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
Code Availability
The code used during the current study is available from the corresponding author on reasonable request.